# LoSA: Long-Short-range Adapter for Scaling End-to-End Temporal Action Localization

### v1

@@ 3.4 Long-Short-range Fusion
The output of the adapters comprises feature representations at different temporal ranges and different intermediate layers of the video backbone. We maintain that these features from intermediate layers should function to enhance the temporal understanding of the last layer features during end-to-end training. We therefore need to strategically fuse these representations to generate TAL-enhanced feature representations for the TAL head that are the most optimal in localizing actions.
@@
For this, after obtaining the short-range and long-range features from each intermediate layer, we introduce a Long-Short-range Fusion module (Fig 2c) to learn to strategically fuse these features along with the last layer features to obtain the TAL-enhanced features that are fed to the TAL head. As shown in Fig 2c, the Long-Short-range Fusion module comprises a short-range gating layer, $\mathsf{Gate}^{\mathsf{sh}}_{i}$, at each intermediate layer $f_{1},\ldots,f_{N-1}$. $\mathsf{Gate}^{\mathsf{sh}}_{i}$ multiplies the temporally concatenated short-range features, obtained from the Short-range Adapter, at layer $f_{i}$, $\mathbf{FS^{X}_{i}}=\{{FS}_{i}^{x_{1}},\ldots,{FS}_{i}^{x_{T}}\}\in\mathbb{R}^{TT_{i}\times C_{i}}$ with a learnable scaling parameter $p^{\mathsf{sh}}_{i}$ to compute the short-range contribution at intermediate layer $f_{i}$ as, $\mathsf{Gate}^{\mathsf{sh}}_{i}(\mathbf{FS^{X}_{i}})=p^{\mathsf{sh}}_{i}\mathbf{FS^{X}_{i}}$. Next, we sum the short-range contributions over all intermediate layers as, $\mathbf{FS^{X}}=\sum_{i=1}^{N-1}\mathsf{Gate}^{\mathsf{sh}}_{i}(\mathbf{FS^{X}_{i}}).$ We note that depending on the network architecture $T_{i}$ may vary in size across intermediate layers. To accommodate that and faciliate fusion, we use linear projection layers, as needed, to transform each $T_{i}\rightarrow T_{N}$, temporal dimension of the last layer $f_{N}$.
@@
Similarly, the Long-Short-range Fusion module comprises a long-range gating layer, $\mathsf{Gate}^{\mathsf{lo}}_{i}$, at each intermediate layer $f_{1},\ldots,f_{N-1}$. $\mathsf{Gate}^{\mathsf{lo}}_{i}$ multiplies the long-range features, obtained from the Long-range Adapter, at layer $f_{i}$, $\mathbf{FL^{X}_{i}}\in\mathbb{R}^{TT_{i}\times C_{i}}~{}\forall~{}i\in\{1\ldots N-1\}$ with a learnable scaling parameter $p^{\mathsf{lo}}_{i}$ to compute the long-range contribution at intermediate layer $f_{i}$ as, $\mathsf{Gate}^{\mathsf{lo}}_{i}(\mathbf{FL^{X}_{i}})=p^{\mathsf{lo}}_{i}\mathbf{FL^{X}_{i}}$. Next, we sum the long-range contributions over all intermediate layers as, $\mathbf{FL^{X}}=\sum_{i=1}^{N-1}\mathsf{Gate}^{\mathsf{lo}}_{i}(\mathbf{FL^{X}_{i}}).$ Similar to short-range gating, we use linear projection layers, as needed, to make temporal dimensions at different layers consistent with that of last layer $f_{N}$, i.e., $T_{N}$.
@@
By doing so, we consider the short-range and long-range contribution of intermediate layers as a residual contribution to the original last layer features. In order to mathematically enforce that, we perform zero initialization on the scaling parameters $\{p^{\mathsf{sp}}_{i}\}^{N-1}_{i=0}$ and $\{p^{\mathsf{tem}}_{i}\}^{N-1}_{i=0}$ of the gating layers $\{\mathsf{Gate}^{\mathsf{sp}}_{i}\}^{N-1}_{i=0}$ and $\{\mathsf{Gate}^{\mathsf{tem}}_{i}\}^{N-1}_{i=0}$ respectively. This ensures that when training starts, the TAL-enhanced features going into the TAL head are effectively the last layer’s original features, providing a stable baseline for leveraging the video backbone $g$. We finally feed $\mathbf{FT^{X}_{N}}$ to the TAL head for generating the localized action snippets as output (Fig 2, bottom-right).
@@ 3.5 Enabling Memory-and-Parameter Efficiency
The unique adapter design of LoSA enables temporal video understanding over the full untrimmed video at each intermediate layer of the video backbone during end-to-end training (Fig 1d). This allows the video backbone, originally trained on trimmed videos, to improve its understanding of the untrimmed video over long-range and short-range temporal context. This is unlike any existing end-to-end TAL method, including both full backbone adaptation and PETL (Fig 1b,c), where there is no mechanism to capture this long-range and short-range temporal understanding of an untrimmed video directly at intermediate layers. A crucial challenge in enabling untrimmed temporal learning at each intermediate layer during end-to-end training is the prohibitively large GPU memory footprint. This is because for TAL, each untrimmed video sample in a training batch involves processing several video clips together. Furthermore, the memory issue is aggravated when leveraging large billion-parameter-plus video backbones. LoSA’s adapter design, performing cross-attention with last layer features, allows the Long-range and Short-range Temporal Adapters to run parallel to the video backbone (as shown in Fig 2). This makes the design memory-and-parameter efficienct and circumvents backpropagating gradients through the video backbone, thereby significantly reducing the GPU memory footprint.
@@ 4 Experiments
Model Backbones. We consider three video backbones: SlowFast-101 [13], VideoMAEv2 (ViT-Base) [34], and VideoMAEv2 (ViT-g) [34]. VideoMAEv2 (ViT-g) has $\sim$1.01 billion parameters and is one of the more recent large video foundation models. It achieves SOTA performance on several video benchmarks including TAL. We select VideoMAEv2 (ViT-g) in order to demonstrate the scaling limitations of existing methods as well as LoSA’s effectiveness in overcoming those limitations. We select SlowFast-101 and VideoMAEv2 (ViT-Base) as these models (along with others of similar sizes) are widely used in the TAL literature. Moreover, it also helps to evaluate the effectiveness of LoSA on different families of model architecture with SlowFast-101 being CNN-based [19] and VideoMAEv2 (ViT-Base) being a Transformer [33]-based model.
Implementation Details. We train LoSA feeding RGB-frames (similar to existing end-to-end TAL methods) as input to the different video backbones with an initial learning rate of 1e-4 for THUMOS-14 and 1e-3 for ActivityNet using a cosine annealing, warmup of 5 epochs, and AdamW [26] optimizer. We use Actionformer as the TAL head with max sequence length of 576 frames at $224\times 224$ spatial resolution with $T^{\prime}=16$ frames. We use temporally-consistent spatial augmentation involving random resizing and cropping and autoaugment [10]. We attach Short-range and Long-range Adapters to all the intermediate layers with $n_{\text{heads}}=4$.
Please refer to the supplementary for additional details.
@@
4.1 Comparison of Head-only vs. End-to-End TAL Training
Table 1 provides a comparison of LoSA with different training strategies available for TAL (as shown in Fig 1). We experiment using THUMOS-14 on SlowFast-101 and VideoMAEv2 (ViT-g) to show the comparison on backbones with sizes at different orders of magnitude. We can observe that for both backbones, LoSA significantly outperforms head-only transfer learning (Fig 1a) by $3.1\%$ and $1.4\%$ respectively on Avg mAP. LoSA also outperforms full backbone adaptation on SlowFast-101 by $1.8\%$ Avg mAP. Full backbone adaption (Fig 1b) results in GPU Out of Memory error (OOM) on VideoMAEv2 (ViT-g) owing to the backbone having 1.02 billion parameters, which prevents even a batch size of one training sample to fit in an A100 GPU.
@@
4.2 Comparison of different Adapter designs for TAL
Since there exists no previous adapter-based PETL method for TAL, we re-purpose some of the existing PETL approaches to work for TAL using publicly available implementations. We consider ST-Adapter, AIM, and UniPT for comparison. For SlowFast-101, LoSA significantly outperforms all existing adapter-based methods by at least $1.4\%$ Avg mAP, highlighting the importance of the design of LoSA’s Long-Short-range Adapter tailored specifically for TAL. On VideoMAEv2 (ViT-g), the original implementation of ST-Adapter and AIM leads to an OOM error because, as Table 2 and Fig 1c show, while their learnable parameter count is significantly less than the full backbone, they are not memory efficient, making them unscalable to billion parameter models like VideoMAEv2 (ViT-g). LoSA works on VideoMAEv2 (ViT-g) because it is both memory and parameter efficient. UniPT has a memory-efficient adapter design, allowing it to operate within the GPU memory limit. Meanwhile, LoSA significantly outperforms UniPT as our adapter design to capture temporal context over different temporal ranges specializes the method for TAL and untrimmed videos.
@@ 4.3 Comparison with State-of-the-Art
As per Table 4.2a, we can observe that LoSA significantly outperforms head-only training (✗ in E2E Adaptation column) on both CNN-based and transformer-based video backbones – SlowFast-101, VideoMAEv2 (ViT-B), and VideoMAEv2 (ViT-g), by $3.1\%$, $1.4\%$, and $1.4\%$ on Avg mAP respectively. Similarly in Table 4.2b, we can see that LoSA outperforms head-only transfer learning on all the video backbones – VideoMAEv2 (ViT-B) and VideoMAEv2 (ViT-g), by $1.3\%$ and $1.5\%$ on Avg mAP respectively. This shows the effectiveness of LoSA in better leveraging and adapting video backbones across different sizes for TAL beyond head-only transfer learning. By doing so, LoSA outperforms all existing TAL methods, including those that use both RGB and optical flow features and those that attempt backbone adaptation, thereby establishing a new SOTA on both THUMOS-14 and ActivityNet-v1.3.
& Gating Strategy Avg mAP Gating w/ random initialization 69.1 Gating w/ one initialization 68.4 Cosine similarity scaling 70.0 Gating w/ zero initialization (Ours) 71.0 Intermediate Layers ($f_{i}$) Avg mAP $f_{1},\ldots,f_{20}$ 44.4 $f_{21},\ldots,f_{39}$ 70.6 $f_{10},\ldots,f_{30}$ 69.4 $f_{30},\ldots,f_{39}$ 69.9 $f_{15},\ldots f_{20}$, $f_{35},\ldots,f_{39}$ 68.9 $f_{1},\ldots,f_{39}$ (All, Ours) 71.0 (a) (b) (c)
@@ 4.4 Ablation
We conduct an ablation study using THUMOS-14 and VideoMAEv2 (ViT-g), as shown in Table 4.3a, to highlight the effectiveness of each integral component of LoSA. The table shows that omitting the Long-range or Short-range Adapter reduces Avg mAP by at least $1.8\%$, indicating the necessity of both in incorporating temporal information at different ranges from the intermediate layers for optimal TAL performance. Next, we conduct an ablation where we remove the Long-Short-range Fusion module and replace it with a simple addition of the features. As Row 3 in Table 4.3a shows, removing the Long-Short-range Fusion module leads to a significant drop of $3.2\%$ in Avg mAP. This shows that along with adapting the long-range and short-range temporal information via the respective adapters, it is also critical to learn how to scale the contribution across the intermediate layers to allow the most relevant long-range and short-range temporal information to be incorporated into the features being fed to the TAL head.
@@ 4.5 Discussion
Gating w/ zero initialization outperforms all other gating strategies. In comparing LoSA’s gating layer strategy with alternatives, our focus is on its unique design and efficiency. As detailed in Sec 3.4, we enforce a zero initialization on the scaling parameter in the long-range and short-range gating layers to enable the long-range and short-range contributions to function as a residual contribution with respect to the last layer features. Table 4.3b shows that this approach yields the highest Avg mAP, outperforming random or one-value initializations. An alternative to parameterized scaling in the gating layer can be to weigh the long-range and short-range contribution of each intermediate layer based on cosine similarity between intermediate features and the last layer features. We find this design also falls short of the gating layer strategy of LoSA. We believe this is because cosine similarity imposes a similarity prior on the scaling parameter, which restricts its ability in scaling the contribution.
@@
Gating parameter learns different values over intermediate layers. Since the scaling parameter in the long-range and short-range gating layers is a learnable parameter, we assess the distribution of the values learned by the scaling parameter post-training across the intermediate layers of the video backbone. For VideoMAEv2 (ViT-g) on THUMOS-14, we find that the range of learnable parameter values to be [-0.01, 0.52] across the long-range and short-range gating in the intermediate layers. This validates two hypotheses. One, all learnable parameters do not collapse trivially to their originally initialized value of 0. At least some of them learn a non-zero value to scale and provide a meaningful residual contribution from the intermediate layers to the TAL-enhanced features, $\mathbf{FT^{X}_{N}}$, going into the TAL head. Two, the learnable parameters exhibit a range of values which indicates that the scaling parameters learns to contribute differently from the intermediate layers as per their importance in improving the TAL-enhanced features, $\mathbf{FT^{X}_{N}}$.
@@
Effect of adapting different intermediate layers. Our experiments with VideoMAEv2 (ViT-g) on THUMOS-14 reveal that attaching Long-range and Short-range Adapters to all 40 transformer layers yields the highest Avg mAP (see Table 4.3c), underscoring the contribution of each layer’s temporal information over different time spans to the TAL head in enhancing localization performance. We can also observe that with the deeper half of layers, $f_{21},\ldots,f_{39}$, we get very close to optimal Avg mAP while shallower half of layers, $f_{1},\ldots,f_{20}$, leads to low Avg mAP. We believe this is due to deeper layers capturing more comprehensive information about the video than shallower layers. While a combination of all layers is optimal, using only the deeper half of layers can suffice when training resources are even more limited.
Figure 3: Qualitative comparison of LoSA with baseline (Head-only) on (a) THUMOS-14 and (b) ActivityNet-v1.3 videos using VideoMAEv2 (ViT-g) showing more accurate action snippets obtained by LoSA.
Visualizations. We conduct a qualitative analysis of the TAL performance on VideoMAEv2 (ViT-g). As shown in Table 1 and 2, only Head-only (Fig 1a) and LoSA (Fig 1d) can work on VideoMAEv2 (ViT-g) without causing GPU OOM error. Fig 3 therefore shows a qualitative comparison of LoSA with head-only transfer learning baseline on THUMOS-14 (Fig 3a) and ActivityNet-v1.3 (Fig 3b).
The figure demonstrates the superior temporal action localization capabilities of LoSA over a conventional head-only transfer learning. The model’s predictions closely match the ground truth, particularly in delineating the nuanced temporal boundaries of the “Pole Vault”, “Hammer Throw”, and “Brushing Teeth” activities. In contrast, the head-only model exhibits clear inaccuracies. These visualizations succinctly demonstrate LoSA’s enhanced precision in TAL with untrimmed videos by introducing an adapter design for end-to-end TAL which can enable temporal understanding over different temporal ranges directly at intermediate layers.

### v2

@@ 3.4 Long-Short-range Gated Fusion
The output of the adapters comprises feature representations at different temporal ranges and different intermediate layers of the video backbone. We maintain that these features from intermediate layers should function to enhance the temporal understanding of the last layer features during end-to-end training. We therefore need to strategically fuse these representations to generate TAL-enhanced feature representations for the TAL head that are the most optimal in localizing actions. For this, after obtaining the short-range and long-range features from each intermediate layer, we introduce a Long-Short-range Gated Fusion module (Fig 2c) to learn to strategically fuse these features along with the last layer features to obtain the TAL-enhanced features that are fed to the TAL head.
@@
As shown in Fig 2c, the Long-Short-range Gated Fusion module comprises a short-range gating layer, $\mathsf{Gate}^{\mathsf{sh}}_{i}$, at each intermediate layer $f_{1},\ldots,f_{N-1}$. $\mathsf{Gate}^{\mathsf{sh}}_{i}$ multiplies the temporally concatenated short-range features, obtained from the Short-range Adapter, at layer $f_{i}$, $\mathbf{FS^{X}_{i}}=\{{FS}_{i}^{x_{1}},\ldots,{FS}_{i}^{x_{T}}\}\in\mathbb{R}^{TT_{i}\times C_{i}}$ with a learnable scaling parameter $p^{\mathsf{sh}}_{i}$ to compute the short-range contribution at intermediate layer $f_{i}$ as, $\mathsf{Gate}^{\mathsf{sh}}_{i}(\mathbf{FS^{X}_{i}})=p^{\mathsf{sh}}_{i}\mathbf{FS^{X}_{i}}$. Next, we sum the short-range contributions over all intermediate layers as, $\mathbf{FS^{X}}=\sum_{i=1}^{N-1}\mathsf{Gate}^{\mathsf{sh}}_{i}(\mathbf{FS^{X}_{i}}).$ We note that depending on the network architecture $T_{i}$ may vary in size across intermediate layers. To accommodate that and faciliate fusion, we use linear projection layers, as needed, to transform each $T_{i}\rightarrow T_{N}$, temporal dimension of the last layer $f_{N}$.
@@
Similarly, the Long-Short-range Gated Fusion module comprises a long-range gating layer, $\mathsf{Gate}^{\mathsf{lo}}_{i}$, at each intermediate layer $f_{1},\ldots,f_{N-1}$. $\mathsf{Gate}^{\mathsf{lo}}_{i}$ multiplies the long-range features, obtained from the Long-range Adapter, at layer $f_{i}$, $\mathbf{FL^{X}_{i}}\in\mathbb{R}^{TT_{i}\times C_{i}}~{}\forall~{}i\in\{1\ldots N-1\}$ with a learnable scaling parameter $p^{\mathsf{lo}}_{i}$ to compute the long-range contribution at intermediate layer $f_{i}$ as, $\mathsf{Gate}^{\mathsf{lo}}_{i}(\mathbf{FL^{X}_{i}})=p^{\mathsf{lo}}_{i}\mathbf{FL^{X}_{i}}$. Next, we sum the long-range contributions over all intermediate layers as, $\mathbf{FL^{X}}=\sum_{i=1}^{N-1}\mathsf{Gate}^{\mathsf{lo}}_{i}(\mathbf{FL^{X}_{i}}).$ Similar to short-range gating, we use linear projection layers, as needed, to make temporal dimensions at different layers consistent with that of last layer $f_{N}$, i.e., $T_{N}$.
@@
By doing so, we consider the short-range and long-range contribution of intermediate layers as a residual contribution to the original last layer features. To mathematically enforce that, we perform zero initialization on the scaling parameters $\{p^{\mathsf{sp}}_{i}\}^{N-1}_{i=0}$ and $\{p^{\mathsf{tem}}_{i}\}^{N-1}_{i=0}$ of gating layers $\{\mathsf{Gate}^{\mathsf{sp}}_{i}\}^{N-1}_{i=0}$ and $\{\mathsf{Gate}^{\mathsf{tem}}_{i}\}^{N-1}_{i=0}$ respectively. This ensures that when training starts, the TAL-enhanced features entering the TAL head are effectively the last layer’s original features, providing a stable baseline for leveraging video backbone $g$. We finally feed $\mathbf{FT^{X}_{N}}$ to TAL head for generating the localized action snippets as output (Fig 2, bottom-right).
@@ 3.5 Enabling Memory-and-Parameter Efficiency
The unique adapter design of LoSA enables temporal video understanding over the full untrimmed video at each intermediate layer of the video backbone during end-to-end training (Fig 1d). This allows the video backbone, originally trained on trimmed videos, to improve its understanding of the untrimmed video over long-range and short-range temporal context. This is unlike any existing end-to-end TAL method, including both full backbone adaptation and PETL (Fig 1b,c), where there is no mechanism to capture this long-range and short-range temporal understanding of an untrimmed video directly at intermediate layers. A crucial challenge in enabling untrimmed temporal learning at each intermediate layer during end-to-end training is the prohibitively large GPU memory footprint. This is because for TAL, each untrimmed video in a training batch involves processing several video clips together. The memory issue is further aggravated when leveraging $>$1 B parameter video backbones. LoSA’s adapter design, performing cross-attention with last layer features, allows the Long-range and Short-range Temporal Adapters to run parallel to the video backbone (Fig 2). This makes the design memory-and-parameter efficient and circumvents backpropagating gradients through the video backbone, thereby significantly reducing the GPU memory footprint.
@@ 4 Experiments
Model Backbones. We consider three video backbones: SlowFast-101 [12], VideoMAEv2 (ViT-Base) [33], and VideoMAEv2 (ViT-g) [33]. VideoMAEv2 (ViT-g) has $\sim$1.01 billion parameters and is among the recent large video foundation models. It achieves SOTA on several video benchmarks including TAL. We select VideoMAEv2 (ViT-g) to demonstrate the scaling limitations of existing methods as well as LoSA’s effectiveness in overcoming those limitations. We select SlowFast-101 and VideoMAEv2 (ViT-Base) as these models are widely used in the TAL literature. This further helps to evaluate the effectiveness of LoSA on different families of model architecture with SlowFast-101 being CNN-based [18] and VideoMAEv2 (ViT-Base) being a Transformer [32]-based model.
Implementation Details. We train LoSA by feeding RGB-frames (similar to existing end-to-end TAL methods) as input to the different video backbones with an initial LR of 1e-4 for THUMOS-14 and 1e-3 for ActivityNet using a cosine annealing, warmup of 5 epochs, and AdamW [25] optimizer. We use Actionformer as the TAL head with max sequence length of 576 frames at $224\times 224$ spatial resolution with $T^{\prime}=16$ frames. We use temporally-consistent spatial augmentation involving random resizing and cropping and autoaugment [10]. We attach Short-range and Long-range Adapters to all the intermediate layers with $n_{\text{heads}}=4$. Please refer to the supplementary for additional details.
4.1 Head-only vs. End-to-End TAL Training
@@
Table 1 provides a comparison of LoSA with different training strategies available for TAL (as shown in Fig 1). We experiment using THUMOS-14 on SlowFast-101 and VideoMAEv2 (ViT-g) to show the comparison on backbones with sizes at different orders of magnitude. We can observe that for both backbones, LoSA significantly outperforms head-only transfer learning (Fig 1a) by $3.1\%$ and $1.4\%$ respectively on Avg mAP. LoSA also outperforms full backbone adaptation on SlowFast-101 by $1.8\%$ Avg mAP. Full backbone adaptation (Fig 1b) results in GPU Out of Memory error (OOM) on VideoMAEv2 (ViT-g) due to the backbone having 1.02 billion parameters, which prevents even a batch of one training sample to fit in an A100 GPU.
4.2 Different Adapter designs for TAL
@@
Since there exists no previous adapter-based PETL method for TAL, we re-purpose some of the existing PETL approaches to work for TAL. We consider ST-Adapter and AIM for comparison. For SlowFast-101, LoSA significantly outperforms all existing adapter-based methods by at least $1.4\%$ Avg mAP, highlighting the importance of the design of LoSA’s Long-Short-range Adapter tailored specifically for TAL. On VideoMAEv2 (ViT-g), the original implementation of ST-Adapter and AIM leads to an OOM error because, as Table 2 and Fig 1c show, while their learnable parameter count is significantly less than the full backbone, they are not memory efficient, making them unscalable to billion parameter models like VideoMAEv2 (ViT-g). LoSA works on VideoMAEv2 (ViT-g) because it is both memory and parameter efficient. Additionally, LoSA’s design, which captures temporal context over different ranges, specializes the method for TAL and untrimmed videos, leading to its significant outperformance.
@@ 4.3 Comparison with State-of-the-Art
As per Table 4.2a, we can observe that LoSA significantly outperforms head-only training (✗ in E2E Adaptation column) on both CNN-based and transformer-based video backbones – SlowFast-101, VideoMAEv2 (ViT-B), and VideoMAEv2 (ViT-g), by $3.1\%$, $1.4\%$, and $1.4\%$ on Avg mAP respectively. Similarly in Table 4.2b, we can see that LoSA outperforms head-only transfer learning on all the video backbones – VideoMAEv2 (ViT-B) and VideoMAEv2 (ViT-g), by $1.3\%$ and $1.5\%$ on Avg mAP respectively. This shows the effectiveness of LoSA in better leveraging and adapting video backbones across different sizes for TAL beyond head-only transfer learning. LoSA outperforms all existing TAL methods, including those that use both RGB and optical flow features and those that attempt backbone adaptation, thereby establishing a new SOTA on both THUMOS-14 and ActivityNet-v1.3.
& Gating Strategy Avg mAP Gating with random initialization 70.0 Gating with one initialization 69.8 Gating with zero initialization (Ours) 71.0 Intermediate Layers ($f_{i}$) Avg mAP $f_{1},\ldots,f_{20}$ 44.4 $f_{21},\ldots,f_{39}$ 70.6 $f_{10},\ldots,f_{30}$ 69.4 $f_{30},\ldots,f_{39}$ 69.9 $f_{15},\ldots f_{20}$, $f_{35},\ldots,f_{39}$ 68.9 $f_{1},\ldots,f_{39}$ (All, Ours) 71.0 (a) (b) (c)
@@ 4.4 Ablation
We conduct an ablation study using THUMOS-14 and VideoMAEv2 (ViT-g), as shown in Table 4.3a, to highlight the effectiveness of each integral component of LoSA. The table shows that omitting the Long-range or Short-range Adapter reduces Avg mAP by at least $0.8\%$, indicating the necessity of both in incorporating temporal information at different ranges from the intermediate layers for optimal TAL performance. Next, we conduct an ablation where we remove the Long-Short-range Gated Fusion module and replace it with a simple addition of the features. As Row 3 in Table 4.3a shows, removing the Long-Short-range Gated Fusion module leads to a significant drop of $1.2\%$ in Avg mAP. This shows that along with adapting the long-range and short-range temporal information via the respective adapters, it is also critical to learn how to scale the contribution across the intermediate layers to allow the most relevant long-range and short-range temporal information to be incorporated into the features being fed to the TAL head.
@@ 4.5 Discussion
Gating with zero initialization outperforms all other gating strategies. In comparing LoSA’s gating layer strategy with alternatives, our focus is on its unique design and efficiency. As detailed in Sec 3.4, we enforce a zero initialization on the scaling parameter in the long-range and short-range gating layers to enable the long-range and short-range contributions to function as a residual contribution with respect to the last layer features. Table 4.3b shows that this approach yields the highest Avg mAP, outperforming random or one-value initializations.
@@
Gating parameter learns different values over intermediate layers. Since the scaling parameter in the long-range and short-range gating layers is a learnable parameter, we assess the distribution of the values learned by the scaling parameter post-training across the intermediate layers of the video backbone. For VideoMAEv2 (ViT-g) on THUMOS-14, we find that the value of learnable parameter ranges in [-0.01, 0.52] across the long-range and short-range gating in the intermediate layers. This validates two hypotheses. One, all learnable parameters do not collapse trivially to their originally initialized value of 0. At least some of them learn a non-zero value to scale and provide a meaningful residual contribution from intermediate layers to the TAL-enhanced features, $\mathbf{FT^{X}_{N}}$, entering the TAL head. Two, the learnable parameters exhibit a range of values which indicates that the scaling parameters learn to contribute differently from the intermediate layers as per their importance in improving $\mathbf{FT^{X}_{N}}$.
@@
Effect of adapting different intermediate layers. Our experiments with VideoMAEv2 (ViT-g) on THUMOS-14 reveal that attaching Long-range and Short-range Adapters to all 40 transformer layers yields the highest Avg mAP (see Table 4.3c), underscoring the contribution of each layer’s temporal information over different time spans to the TAL head in enhancing localization performance. We can also observe that with the deeper half layers, $f_{21},\ldots,f_{39}$, we get very close to optimal Avg mAP while shallower half layers, $f_{1},\ldots,f_{20}$, leads to low Avg mAP. We believe this is due to deeper layers capturing more comprehensive information about the video than shallower layers. While a combination of all layers is optimal, using only the deeper half layers can suffice when training resources are limited.
Significant gains on ActivityNet-v1.3. Tab. 4.2 shows that improvements between successive works in recent years on ActivitNet-v1.3 is generally around $0.5-1\%$ Avg mAP. In contrast, LoSA can improve by more than 1% Avg mAP compared to any existing baseline, showing the significant performance gain achieved by LoSA on ActivityNet-v1.3.
Figure 3: Sensitivity analysis on THUMOS-14 using [2]. mAPN denotes normalized mAP at tIoU=0.5 with N average ground truth segments per class. Top: LoSA w/o Long-Short-Adapter. Bottom: LoSA (Ours). Performance for both XS and XL improves significantly with our method LoSA (bottom) compared to the baseline, LoSA w/o Long-Short-range Adapter (top).
LoSA improves both long- and short-action instances. Fig 3 shows a sensitivity analysis on THUMOS-14 using VideoMAEv2 (VIT-g), taking into account coverage, length, and number of action instances. We show a comparison between LoSA w/o Long-Short-range Adapter (baseline, the top row) and LoSA (ours, the bottom row). Refer to supplementary for setup details. We observe that the performance for both XS and XL improves significantly with our method LoSA compared to the baseline, LoSA w/o Long-Short-range Adapter. We believe this is because LoSA enables untrimmed temporal learning at different temporal ranges in the intermediate layers via the Long- and Short-range Temporal Adapters. This enables capturing complex scene details as well as fine-grained information required for long- and short- duration action instances respectively.